A particle P of mass 2m lies on a rough horizontal table. P is connected by a light inextensible string passing over a smooth pulley fixed at the edge of the table to another particle Q of mass 5m, hanging freely. The system is released from rest with the string taught and the hanging part vertical. If the acceleration of the system is of magnitude 4g/7, find
(a) the tension in the string
(b) the coefficient of friction between the first particle and the table.
1
Expert's answer
2021-02-06T10:07:22-0500

Let's solve the problem.

we have "a = \large\frac{4g}{7}"

first of all we mark the directions of x and y forces which impact to P and Q particles


 for P particle: "N = 2mg \space\space\space\space 2ma = T - F_{ff} = T - \mu N = T - \mu 2mg"

for Q particle: "5ma = 5mg - T"

"\begin{cases}
 2ma = T - \mu 2mg \\
 5ma = 5mg - T
\end{cases}"    "+ \to 7ma = 5mg-2\mu mg / : m"

                                                                  "7a = 5g - 2\mu g"

                                                                "\mu = \large\frac{5g-7a}{2g}" "= 2.5-3.5\large\frac{a}{g}" = "\large\frac{5}{2}-\frac{7}{2}*\frac{4}{7}" "= 0.5"

(b) the coefficient of friction between the first particle and the table is 0.5

(a)  the tension in the string is "T = 5m(a+g) = 5m(g + \large\frac{4g}{7}) = \frac{55mg}{7}"